OpenStudy (anonymous):

write an equation in the form y=mx+b for the line that passes through the point (-5,-1) and has a slope of -2

pooja195 (pooja195):

\[\huge~y-y_1=m(x-x_1)\]

pooja195 (pooja195):

m=slope x1,y1

OpenStudy (anonymous):

-6

pooja195 (pooja195):

-6 is what?

pooja195 (pooja195):

\[\huge~y+1=-2(x+5)\]

pooja195 (pooja195):

-2 * x = -2 -2 * 5 = ?

OpenStudy (anonymous):

-10

pooja195 (pooja195):

\[\huge~y+1=-2x-10\] -10+1=?

OpenStudy (anonymous):

-9

pooja195 (pooja195):

y=-2x-9 <---final : )

OpenStudy (anonymous):

thank you for everything again really appreciate it

pooja195 (pooja195):

:)